### v2

# How Cohesive Are Community Search Results on Online Social Networks?: An Experimental Evaluation

Abstract.
Recently, numerous community search methods for large graphs have been proposed, at the core of which is defining and measuring cohesion. This paper experimentally evaluates the effectiveness of these community search algorithms w.r.t. cohesiveness in the context of online social networks. Social communities are formed and developed under the influence of group cohesion theory, which has been extensively studied in social psychology. However, current generic methods typically measure cohesiveness using structural or attribute-based approaches and overlook domain-specific concepts such as group cohesion. We introduce five novel psychology-informed cohesiveness measures, based on the concept of group cohesion from social psychology, and propose a novel framework called CHASE for evaluating eight representative CS algorithms w.r.t. these measures on online social networks. Our analysis reveals that there is no clear correlation between structural and psychological cohesiveness, and no algorithm effectively identifies psychologically cohesive communities in online social networks. This study provides new insights that could guide the development of future community search methods.
copyright: noneconference: xx; xx; xx
1. Introduction
Real-world networks frequently exhibit community structures (Girvan and Newman, 2002), making community search (CS) a key area of focus in graph search (Fang et al., 2020). It seeks to retrieve cohesive subgraphs from an input graph based on query nodes. Communities extracted by CS algorithms have potential applications in event organization, recommendations, protein complex identification, and targeted advertising  (Fang et al., 2020; Gao et al., 2021). In particular, human beings thrive in communities, forming connections through shared experiences and values, which promote cooperation, resilience, and a sense of belonging (Baumeister and Leary, 2017; Wellman, 2001). Thus, community search is particularly important for applications involving human interaction, such as those in online social networks.
Extensive research on the community search problem has primarily focused on enhancing the efficiency, scalability, and quality of results (Fang et al., 2020). Central to this issue is the concept of cohesion within communities. Most CS algorithms define communities through subgraph cohesiveness metrics derived from structural network information, such as k-core, k-truss, and k-ECC, along with attribute data (Fang et al., 2020). These measures are generic, meaning they are not tied to any specific application domain. Recently, several studies have investigated learning-based methods for identifying cohesive subgraphs, which integrate both structural information and vertex attributes into neural network models, thereby removing the need for predefined rules for communities (Gao et al., 2021; Li et al., 2023; Hashemi et al., 2023; Fang et al., 2024; Wang et al., 2024a, b). In this paper, we evaluate the cohesiveness of communities produced by these existing techniques within online social networks.
In online social networks, communities are fundamentally characterized as groups of individuals connected through “invisible bonds”, a phenomenon that has been widely studied in social psychology as cohesion or group cohesion since the 1950s (Festinger, 1950; Zimbardo et al., 2003). Most psychology studies recognize cohesion as multi-dimensional (Salas et al., 2015; McLeod and Von Treuer, 2013; Carron and Brawley, 2000) and have demonstrated its validity in various contexts, such as sports (Carron, 1982; Pescosolido and Saavedra, 2012), work (Lu, 2015; Paul et al., 2016), and technology-mediated learning environments (Zamecnik et al., 2022), as well as in both offline and online settings (Galyon et al., 2016; Lyles et al., 2018; Cai, 2023). Consequently, algorithmically identified communities can be hypothesized to manifest psychologically grounded cohesion properties. Notably, structural cohesiveness metrics used in the CS problem have their origins in social psychology (Wasserman and Faust, 1994; Fortunato, 2010).
Existing CS methods evaluate their solutions in three main ways (Das et al., 2021; Li et al., 2017; Miao et al., 2022; Lin et al., 2024; Wang et al., 2024b, a; Zhang et al., 2023): (a) By comparing retrieved communities to known ground-truth communities using various performance metrics, though the interpretation and definition of ground truth vary (Yang and Leskovec, 2012; Yin et al., 2017; Leskovec et al., 2007; Madani et al., 2022; Klymko et al., 2014) (b) By assessing communities against those produced by other algorithms through “effectiveness metrics” or “goodness metrics” that measure community characteristics (Behrouz et al., 2022; Lin et al., 2024), especially when ground-truth is unavailable. (c) By conducting case studies to demonstrate the interpretability and significance of the communities identified by CS algorithms.
None of the existing evaluation approaches systematically assesses the effectiveness of cohesiveness measures in networks involving interacting individuals. Consequently, they fail to address key questions specific to online social networks, such as: (1) Do the generic cohesiveness measures from different CS techniques accurately capture group cohesion in online social networks? (2) How cohesive are the communities identified by these techniques when evaluated using measures based on group cohesion in social psychology?
In this paper, we address these key questions about the cohesiveness of communities identified by community search (CS) algorithms in online social networks. We begin by discussing how research on group cohesion can inform the assessment of cohesion in online social networks. We critique the limitations of existing generic cohesiveness measures used by various CS techniques in capturing the multi-dimensional nature of group cohesion and introduce a set of psychology-informed cohesiveness measures grounded in social psychology’s definition of cohesion to address them. Additionally, we present a framework called CHASE to conduct extensive experiments comparing the cohesiveness of results from eight representative CS techniques using real-world social network data.
Our experimental evaluation reveals three key findings. First, existing algorithms, when using the same query node, often identify communities that differ significantly in terms of structural and psychology-based cohesiveness, with varying query hit rates. Second, recent learning-based algorithms tend to produce communities with low structural cohesiveness or fail to identify valid communities. Third, there is no clear relationship between structural cohesiveness and psychology-based cohesion. More importantly, no current algorithm identifies cohesive communities in online social networks that align with the concept of cohesion in social psychology.
In summary, this work makes the following novel contributions. (a) It is the first study to assess the cohesiveness of social network communities generated by various community search methods, analyzed through the lens of cohesion theories in social psychology. (b) It introduces a set of psychology-informed cohesiveness measures for evaluation, which is unavailable in existing community search evaluation toolkits. (c) The experimental findings presented in this work are novel and have not been reported in previous studies.
2. Related Work
Fang et al. (2020) evaluated CS results from traditional algorithms on simple and attributed graphs. For simple graphs, it experimentally assesses the average running times and reports community quality using four generic structure-based measures: diameter, degree, density, and clustering coefficient. For attributed graphs, the evaluation focuses on keyword-based CS solutions for efficiency and F1 scores. In social psychology, cohesion measures are developed based on a specific definition and are tailored to a target population. Their quality is assessed through reliability and validity (Bhattacherjee, 2012), focusing on how well these measures align with theoretical constructs (Shaver, 2015; Blackstone, 2018). Recently, Bhowmick et al. (S Bhowmick et al., 2024) introduced the paradigm of psychology-informed design and reviewed existing social influence computation techniques that embrace it. None of these efforts undertakes a systematic evaluation of the cohesiveness of CS solutions on online social networks through the lens of social psychology.
3. The Community Search Problem
In this section, we review the community search problem and describe a taxonomy of existing CS algorithms. We begin by introducing some terminology to facilitate the exposition of the paper.
Terminology. An online social network can be modeled as a directed multigraph $\mathcal{G}=(\mathcal{U},\mathcal{E})$, where $\mathcal{U}$ indicates the set of users (vertices) and $\mathcal{E}$ includes edges representing individual and interaction activities. Individual activity occurs when a user $u_{i}$ posts information at time $t$ without targeting anyone, represented as a self-loop $e_{ii}^{t}\in\mathcal{E}$. Interaction activity refers to an interaction between user $u_{i}$ and $u_{j}$ at time $t$, forming a directed edge $e_{ij}^{t}\in\mathcal{E}$. An induced subgraph $H=(\mathcal{U}_{H},\mathcal{E}_{H})$ of $\mathcal{G}$ consists of a vertex subset from $\mathcal{G}$ and all edges between them.
Problem Definition. Based on (Sozio and Gionis, 2010; Fang et al., 2020), community search problem can be formalized as follows:
Problem 3.1.
[Community Search (CS) Problem] Given a graph $G=(V,E)$ and a set of query nodes $Q\subseteq V$, the community search problem seeks to find a community that contains $Q$ and satisfies the following properties:
  • (i)
    Connectivity: Vertices in the community are connected;
  • (ii)
    Cohesiveness: Vertices in the community are intensively linked to each other based on a particular goodness metric.
Note that $\mathcal{G}$ can be reduced to a simple graph $G=(V,E)$ where $V=\mathcal{U}$ and each set of parallel edges in $\mathcal{E}$ between vertices can be reduced to undirected or directed connections.
Taxonomy of CS Algorithms. Following (Wang et al., 2024a), we classify existing CS works into traditional and learning-based categories.
Traditional CS algorithms. Generally, traditional CS algorithms formulate a community by specifying its subgraph properties (Fang et al., 2020). The commonly adopted cohesiveness property is structural-based, including k-core (Sozio and Gionis, 2010; Cui et al., 2014; Barbieri et al., 2015; Yao and Chang, 2021), k-truss (Huang et al., 2014; Akbas and Zhao, 2017; Liu et al., 2020; Zhang et al., 2023), k-clique (Yuan et al., 2017), and k-ECC (Chang et al., 2015; Hu et al., 2016, 2017). These metrics are often referred to as structural cohesiveness metrics (structural metrics for brevity). Additionally, studies incorporate attribute information to formalize cohesiveness by refining existing structural metrics or proposing new attribute cohesiveness metrics alongside the structural ones. Examples of the former approach include attributed $pkd$-truss community (Xie et al., 2021), $(k,\theta)$-community (Miao et al., 2022), $(k,p)$-core (Lu et al., 2022), and $\beta$-temporal core (Lin et al., 2024). Examples of the latter include keyword cohesiveness (Fang et al., 2016), spatial cohesiveness (Fang et al., 2017; Chen et al., 2018), and query cohesiveness (Das et al., 2021).
Learning-based CS algorithms. Due to the structural inflexibility of traditional CS algorithms, recently learning-based methods have been proposed to retrieve communities by learning their hidden patterns without relying on predefined structures. Depending on their reliance on ground-truth communities for training, they can be further categorized as supervised (Jiang et al., 2022; Chen et al., 2023; Hashemi et al., 2023; Fang et al., 2024; Wang et al., 2024a), semi-supervised (Li et al., 2023), and unsupervised (Wang et al., 2024b). However, most learning-based solutions weakly correlate with community cohesiveness, struggle with large graphs due to neural network constraints, and risk of model invalidation when the graph topology changes (Wang et al., 2024a).
4. Cohesiveness Framework for Social Networks
This section outlines the methodology for measuring cohesion based on social psychology and highlights the limitations of current generic cohesiveness measures used in community search techniques. The goal is to determine whether these generic measures align with social psychology’s concept of cohesion, which is crucial for effective community retrieval in online social networks.
4.1. Methodology
In social psychology, measurements are developed from proposed concepts or adapted from existing ones to fit different study subjects and contexts (Shaughnessy et al., 2000; Blackstone, 2018). With well-studied cohesion concepts and measures available, in Figure 1, we summarize a workflow to adapt existing cohesiveness measures in social psychology to the context of social networks. Here, “measures” particularly refer to questionnaires, which are widely used in psychological studies to assess cohesion (Santoro et al., 2015; Forsyth, 2021; Zamecnik et al., 2022).
We start by selecting a cohesion concept from psychological studies and accessing its measures. To ensure consistent positive scoring of cohesiveness, reversed-word items, introduced alongside positively-worded items to reduce response style bias (Suárez Álvarez et al., 2018), are corrected. Besides, two common adaptation methods used in group cohesion studies—rewording and removing inapplicable items (Spink and Carron, 1994; Bollen and Hoyle, 1990)—are applied to fit the remaining items into our study context.
Cohesion Concept and Measures. In this study, we select a widely accepted definition proposed by Carron (1982), which defines cohesion as “a dynamic process which is reflected in the tendency for a group to stick together and remain united in the pursuit of its goals and objectives”. It recognizes cohesion as multidimensional and dynamic that changes its extent and forms throughout group formation, development, maintenance, and dissolution (Carron and Brawley, 2000).
To measure this concept, a conceptual model has been proposed that identifies two key perceptions binding members to their group: a member’s perception of the group as a whole (labeled as group integration (GI)) and a member’s personal attraction to the group (labeled as individual attractions to group (ATG)). Each perception is further divided into task and social aspects, resulting in four correlated constructs that influence group cohesion: group integration-task (GI-T), group integration-social (GI-S), individual attractions to group-task (ATG-T), and individual attractions to group-social (ATG-S) (Carron et al., 1985). The Group Environment Questionnaire (GEQ) derived from this model consists of 18 items and has demonstrated its reliability and validity in various contexts (Carron et al., 1985; Blanchard et al., 2000; Anderson and Warner, 2017). The cohesion score for a group is calculated by averaging individual GEQ responses (Carron et al., 2003).
Measure Adaptation. Regarding our study context, we focus on the social scales of the GEQ, which emphasize social-oriented motivations driving user interactions  (Syn and Oh, 2015; Dindar and Dulkadir Yaman, 2018; Alsalem, 2019; Shin, 2010). Following our workflow, the GEQ is adapted into Table 1, which includes four items for the ATG-S scale and two items for the GI-S scale.
Figure 1. Workflow for adapting cohesiveness measures.
4.2. Comparison
Table 2 summarizes whether existing measures (introduced in Section 3) capture the cohesion features outlined in the adapted GEQ (Table 1). Observe that both ATG and GI items incorporate underlying structural information, yet in distinct ways. Moreover, ATG places greater emphasis on personal opinions, carrying more psychological insights that influence a user’s membership. Items are labeled “S” for structural information and “P” for psychological information. $\checkmark$”, $\triangle$”, and “$\times$” denote that an item is fully captured, partially captured, or not captured by existing measures, respectively.
Intuitively, all existing structural cohesiveness measures, which focus on the subgraph interconnections, successfully capture Item 13. While strong topological connections may suggest users’ enjoyment, friendship formation, and group engagement (Items 1, 5, 7, and 9), these psychological elements are not explicitly characterized. Similarly, it remains unclear to what extent keywords cohesiveness, which captures users’ attribute similarity, and spatial cohesiveness, which reflects spatial relationships, accurately represent ATG-S items. While query cohesiveness incorporates network information to characterize the user’s temporal activeness, its representation of psychology-based cohesiveness is still implicit and insufficient.
Notably, none of the existing metrics capture users’ willingness to interact within a group (Item 11). Furthermore, they cannot be combined to capture all items in Table 1.
Table 1. Adapted GEQ for online social networks.
Scale Item
ATG-S 1. I enjoy being a part of the social activities of this group.
5. Some of my best friends are in this group.
7. I enjoy interactions in my group more than in other groups.
9. For me, this group is one of the most important social groups to which I belong.
GI-S 11. Members of our group would rather interact with each other than post information individually.
13. Our group members frequently interact with each other.
Table 2. Comparison of existing CS cohesiveness metrics and social psychology-based cohesion.
Item (Info) $k$-core $k$-truss $k$-clique $k$-ECC Attributed $pkd$-truss community $(k,\theta)$-community $(k,p)$-core $\beta$-temporal core Keyword cohesiveness Spatial cohesiveness Query cohesiveness
1 (S & P) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
5 (S & P) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
7 (S & P) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
9 (S & P) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
11 (S) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
13 (S) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
5. Psychology-informed Cohesiveness Measures
The preceding section highlighted the limitations of the generic cohesiveness measures in capturing psychological cohesion in social networks. To further assess and compare their effectiveness, in this section, we propose a set of psychology-informed measures that mathematically quantifies group cohesion indicators described in items within Table 1.
Intuitively, for ATG-S items, our formulation focuses on the relationship between individual users and the community, yielding individual scores, which are then averaged into a community score. For GI-S items, the community is treated as a single unit to report a group-level score. All measures are positively scored.
5.1. Measures for ATG-S Items
Enjoyment Index (EI). Item 1 emphasizes the user’s enjoyment level during social interactions. Enjoyment is often measured concerning positive emotions or needs fulfillment during interactions (Syn and Oh, 2015; Kawamichi et al., 2016; Lin et al., 2009). Accordingly, we identify the enjoyment index as a cohesiveness indicator, quantified by the cumulative sentiments elicited during interactions.
Empirical findings indicate that an individual’s sentiment is shaped by the sentiments and frequency of prior interactions (Kawamichi et al., 2016; Bareket-Bojmel and Shahar, 2011). Individuals prioritize emotion-congruent stimuli and respond more quickly to emotional stimuli than neutral ones (Niedenthal and Setterlund, 1994; Vinson et al., 2014; Yang et al., 2014; Kissler et al., 2009; Garcia et al., 2016). Therefore, we focus on the sentiment polarity of each interaction (positive, neutral, or negative, represented as 1, 0, or -1). We model the impact of sentiments as a sentiment-aware excitation function, which quantifies the excitation degree of previous interactions for the sentiment of the current interaction. Then the elicited sentiment can be modeled as the sentiment conveyed in the interaction scaled by the excitation degree.
The sentiment-aware excitation function is constructed based on a basic excitation function, represented as a linear combination of the basic decay functions $\Phi(\cdot)$ (Rahimi et al., 2023) while introducing a sign function to determine whether past activities $\mathcal{E}^{\prec t}$ enhance or diminish the sentiment of the interaction $e^{t}$ at time $t$ ($t^{\prime}<t$), as shown below.
$\lambda(t,e^{t},\mathcal{E}^{\prec t})=\lambda_{0}+\sum_{e^{t^{\prime}}\in\mathcal{E}^{\prec t}}sgn(Senti(e^{t^{\prime}}),Senti(e^{t}))\cdot\Phi(t-t^{\prime})$
Wherein, $\lambda_{0}$ represents the basic sentiment perception level elicited by a single interaction. We set $\lambda_{0}=1$ to ensure that the elicited sentiment matches the original sentiment in the absence of prior activities. The function $Senti(\cdot)\in[-1,1]$ extracts the sentiment from activities, and the sign function is defined as $sgn(Senti(e^{t^{\prime}}),Senti(e^{t}))=Senti(e^{t^{\prime}})\cdot Senti(e^{t})$. The sentiment extraction methods and decay function selection will be introduced in Section 6. Formally, the elicited sentiment is defined as below.
Definition 5.1.
[Elicited Sentiment] Given an interaction activity $e^{t}$ at time $t$, the elicited sentiment is defined as the interaction sentiment scaled by the sentiment excitation degree as follows:
$ESenti(t,e^{t},\mathcal{E}^{\prec t})=Senti(e^{t})\cdot\lambda(t,e^{t},\mathcal{E}^{\prec t}),$
where $\lambda=\max(\lambda,0)$ to ensure non-negative excitation.
Let $\mathcal{E}_{H,i\cdot}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},j\neq i,t\leq t_{cur}\}$ be activities where user $u_{i}$ interacts with other group members up to time $t_{cur}$. Assuming that more recent activities better capture user interests (Sarker et al., 2019), we weight them by a decay function $\Phi(\cdot)$ and define the enjoyment index as follows.
Definition 5.2.
[Enjoyment Index (EI)]Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and a user $u_{i}\in H$, $u_{i}$’s EI is defined as the cumulative elicited sentiments generated during interactions between $u_{i}$ and others in $H$ by $t_{cur}$:
$\mathbb{EI}_{i}(\mathcal{G},H,t_{cur})=\sum_{e^{t}\in\mathcal{E}_{H,i\cdot}^{t_{cur}}}ESenti(t,e^{t},\mathcal{E}_{H,i\cdot}^{\prec t})\cdot\Phi(t_{cur}-t),\$
where $\mathcal{E}_{H,i\cdot}^{\prec t}=\{e_{xy}^{t^{\prime}}\mid e_{xy}^{t^{\prime}}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},j\neq i,t^{\prime}<t\}$.
Sentimental Interaction Tendency (SIT). Now consider Item 5 in Table 1. Having best friends in the group indicates cohesiveness, prompting the need to measure friendships among group members. Friendship generally develops through the exchange of information (Chan and Lo, 2014) and is characterized by reciprocity, closeness, and emotional attachment (Amichai-Hamburger et al., 2013; Paxton and Moody, 2003). Compared to ordinary friends, best friends share greater intimacy, reciprocity, and stability (Berndt, 1982; Roberto and Kimboko, 1989; Weisz and Wood, 2005). For simplicity, we propose sentimental interaction tendency to characterize friendships developed through cumulative information exchange, with higher values denoting closer ties.
Let $\mathcal{M}_{H,i}^{t_{cur}}=\{u_{j}\mid u_{j}\in\mathcal{U}_{H},j\neq i,\exists t_{1},t_{2}\leq t_{cur}:e_{ij}^{t_{1}},e_{ji}^{t_{2}}\in\mathcal{E}_{H}\}$ denotes the set of subgroup users that mutually interact with $u_{i}$. Then, their interactions are characterized as $\mathcal{E}_{H,\{i,j\}}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},t\leq t_{cur}\}$. The cumulative, time-decaying sentiments in these interactions are computed as follows.
$Accu_{ij}(\mathcal{G},H,t_{cur})=\sum_{e^{t}\in\mathcal{E}_{H,\{i,j\}}^{t_{cur}}}ESenti(t,e^{t},\mathcal{E}_{H,\{i,j\}}^{\prec t})\cdot\Phi(t_{cur}-t)$
Note that $Accu_{ij}(\cdot)$ is symmetric, i.e., $Accu_{ij}(\mathcal{G},H,t_{cur})=\linebreak Accu_{ji}(\mathcal{G},H,t_{cur})$, reflecting the reciprocity of friendship.
Definition 5.3.
[Sentimental Interaction Tendency (SIT)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and a user $u_{i}\in H$, $u_{i}$’s SIT is defined as the accumulated elicited sentiments between $u_{i}$ and each user with whom they have mutual interactions up to $t_{cur}$:
$\mathbb{SIT}_{i}(\mathcal{G},H,t_{cur})=\sum_{u_{j}\in\mathcal{M}_{H,i}^{t_{cur}}}Accu_{ij}(\mathcal{G},H,t_{cur}).$
Comparative Enjoyment Degree (CED). Users often engage in multiple groups in online social networks (Tan and Lee, 2015). They perceive greater cohesiveness in a specific group when it positively contributes to their enjoyment, as indicated by Items 7 and 9. This leads us to propose comparative enjoyment degree to measure how much more enjoyment a user gets from community interactions than those outside the community.
Definition 5.4.
[Comparative Enjoyment Degree (CED)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and a user $u_{i}\in H$, the CED of $u_{i}$ at time $t_{cur}$ is defined as the difference between the enjoyment derived from interactions within and outside $H$ by $t_{cur}$ :
$\begin{split}\mathbb{CED}_{i}(\mathcal{G},H,t_{cur})=&\sum_{e^{t}\in\mathcal{E}_{H,i\cdot}^{t_{cur}}}ESenti(t,e^{t},\mathcal{E}_{H,i\cdot}^{\prec t})\cdot\Phi(t_{cur}-t)\\ &-\sum_{e^{t}\in\mathcal{E}_{\mathcal{G}\setminus H,i\cdot}^{t_{cur}}}ESenti(t,e^{t},\mathcal{E}_{\mathcal{G}\setminus H,i\cdot}^{\prec t})\cdot\Phi(t_{cur}-t),\end{split}$
where $\mathcal{E}_{\mathcal{G}\setminus H,i\cdot}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{\mathcal{G}},\{x,y\}=\{i,j\},u_{j}\notin\mathcal{U}_{H},t\leq t_{cur}\}$ represents $u_{i}$’s interactions with users outside $H$.
5.2. Measures for GI-S Items
Group Interaction Preference (GIP). Item 11 captures the group’s preference for group interaction over individual posting, which can be formulated as group interaction preference (GIP). Let $\mathcal{E}_{H,inter}^{t_{cur}}=\{e_{ij}^{t}\mid e_{ij}^{t}\in\mathcal{E}_{H},j\neq i,t\leq t_{cur}\}$ and $\mathcal{E}_{H}^{t_{cur}}=\{e_{ij}^{t}\mid e_{ij}^{t}\in\mathcal{E}_{H},t\leq t_{cur}\}$ denote the set of interaction activities and all activities in $H$ up to $t_{cur}$, respectively, where $\mathcal{E}_{H,inter}^{t_{cur}}\subseteq\mathcal{E}_{H}^{t_{cur}}$. The GIP ranges from 0 to 1 and is defined as follows.
Definition 5.5.
[Group Interaction Preference (GIP)] Given a graph $\mathcal{G}$, community $H$ and current time $t_{cur}$, GIP is defined as the ratio of group members’ interactions to their overall activities:
$\mathbb{GIP}(\mathcal{G},H,t_{cur})=\frac{|\mathcal{E}_{H,inter}^{t_{cur}}|}{|\mathcal{E}_{H}^{t_{cur}}|}.$
Group Interaction Density (GID). Lastly, Item 13 emphasizes the interplay between member relationships and interaction frequency. We quantify it by extending the weighted density (Liu et al., 2009) with a temporal normalization. Let the observation window be $W_{t_{cur}}=(t_{cur}-t_{0})/\tau_{t}$, where $t_{0}$ is the start of observation and $\tau_{t}$ is the time-unit (e.g., day, week, month). We call the resulting metric the group interaction density (GID).
Definition 5.6.
[Group Interaction Density (GID)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and observation window $W_{t_{cur}}$, GID measures the interaction frequency per possible user pair per unit time within $H$ by $t_{cur}$:
$\mathbb{GID}(\mathcal{G},H,t_{cur})=\frac{|\mathcal{E}_{H,inter}^{t_{cur}}|}{|\mathcal{U}_{H}|\cdot(|\mathcal{U}_{H}|-1)\cdot W_{t_{cur}}}.$
Remark. Note that these five measures for assessing cohesion can be merged into a single cohesiveness score through an aggregated linear combination of weights. However, our work does not aim to calculate such a score. Instead, the evaluation shall focus on analyzing the quality of results produced by existing CS techniques w.r.t. each of these measures, enabling us to get a fine-grained understanding of how they perform against each item in Table 1.
6. Experimental Settings
In this section, we introduce the datasets, representative community search algorithms, evaluation framework, and parameter settings used to set up the experiments. The codebase is available at (Zhao, 2025).
6.1. Dataset
Due to the lack of sentiment data in most publicly available datasets (e.g., SNAP (Leskovec and Krevl, 2014)), we constructed four real-world datasets from X (formerly Twitter) to facilitate evaluation. Table 3 summarizes their topological characteristics, where $|\mathcal{U}|$, $|\mathcal{E}|$, and $|\mathcal{T}|$ denote the number of vertices, edges, and edge timestamps, respectively. $Deg_{avg}$ represents the average node degree, with self-loops and parallel edges counted.
Table 3. Dataset statistics.
Dataset $|\mathcal{U}|$ $||\mathcal{E}|$$|\mathcal{T}|$ Density $Deg_{avg}$
BTW17 7,721 12,839 11,890 0.0002 3.33
Chicago_COVID 4,917 246,633 246,001 0.0102 100.32
Crawled_Dataset144 10,388 119,836 116,639 0.0011 23.07
Crawled_Dataset26 21,263 108,538 104,575 0.0002 10.21
Data Collection. The BTW17 dataset contains tweets related to the 19th German election campaigns, with over 1.2 million tweets from active users (Kratzke, 2017). The Chicago_COVID dataset contains tweets related to the COVID-19 pandemic and geolocated near Chicago (Gordon, 2023). Due to its large volume, we sampled 125k tweets for this study. Besides, we collected Crawled_Dataset144 and Crawled_Dataset26 datasets in December 2022 using the official API (Platform, 2022). Starting with single users randomly selected from real-time tweets, we employed snowball sampling to gather the users’ original tweets and associated replies (conversation utterances). Newly identified users were recorded for future data collection. This process resulted in over 200k original tweets for each dataset. For all four datasets, we extracted metadata for each user, including user_id. The tweet metadata includes conversation_id, tweet_id, user_id, in_reply_to_user_id, tweet_text, and timestamp. The number of users and tweet data elements varied across datasets. We retained only English tweets, removed retweets, quoted tweets, duplicates, and empty entries, and cleaned the tweet text by removing newlines and URLs.
Figure 2. Prompt for sentiment analysis.
Sentiment Analysis. The preprocessed tweets were then subjected to a sentiment analysis task to determine their sentiment scores. Following common practice, we treat it as a three-class classification problem (Chang et al., 2024). To ensure superior analysis results, we adopted the Llama3-8B, an open-source Large Language Model (LLM) from Meta (AI@Meta, 2024), and used a prompt-based approach. The prompt was adapted from (Kheiri and Karimi, 2023) to assign sentiment labels (see Figure 2), which were mapped to 1, 0, and -1, respectively. Both top_p and temperature were set to 0.01 for more deterministic responses. Only the tweets successfully assigned labels were retained.
Network Construction. Networks were constructed based on the reply relationships (i.e., who replies to whom), including interaction and self-posting tweets (i.e., initial posts not directed at specific users). Isolated nodes were removed. Since the constructed networks consist of components with their size distribution highly skewed, only the largest component in each network was considered, resulting in the networks in Table 3.
6.2. Representative CS Algorithms
We selected representative CS algorithms based on the following criteria: (a) Their codebases are publicly available; (b) They cover the two categories introduced in Section 3; (c) They consider various information to characterize cohesiveness, potentially capturing different items in the adapted GEQ (Table 2). As a result, we choose three $k$-core-based algorithms (ALS, WCF-CRC, CSD), three $k$-truss-based algorithms (ST-Exa, Repeeling+, I2ACSM), and two unsupervised learning algorithms (TransZero-LS, TransZero-GS), as summarized in Table 4. Note that although (Das et al., 2021) could be a useful technique for our study (based on Table 2), unfortunately, its codebase is not available. The last column in Table 4 indicates whether the algorithms use ground-truth for evaluation, with most of them not using it. Note that the ground-truth adopted in the three approaches is not based on group cohesion theory and lacks data elements (e.g., temporality, sentiments) to calculate our proposed measures. Therefore, we cannot exploit them for our evaluation. We now briefly describe these algorithms.
ALS identifies the $\beta$-temporal proximity core with the largest $\beta$ in a temporal graph (Lin et al., 2024), integrating temporal proximity with structural cohesiveness seamlessly. It requires specifying the teleportation probability $\alpha$ to compute temporal proximity. WCF-CRC searches on dynamic networks, represented as a sequence of time-variant weighted graph instances ($\mathbb{G}={G_{t_{1}},G_{t_{2}},\ldots,G_{t_{T}}}$) (Tang et al., 2022). It aims to find a Time Interval based ($\theta,k$)-core Reliable Community (CRC) with the highest reliability score, where a ($\theta,k$)-core in a $G_{t}$ is a $k$-core where all edge weights are at least $\theta$. CSD finds a $(k,l)$-core in the directed graph, where $k$ and $l$ denote the minimum in-degree and out-degree of subgraph nodes (Chen et al., 2021).
ST-Exa identifies a subgraph with the largest min-support based on the query node, while satisfying the size constraint $[l,h]$. It is parameter-free, eliminating the need to specify the cohesiveness constraint. Repeeling+ is a peeling-based algorithm for finding a maximal $D$-truss containing the query node in the snapshot digraph $G^{t}$ (Liao et al., 2024). $G^{t}$ is induced from a streaming directed graph using a sliding window length $\tau$ and a stride $\beta$. The algorithm’s parameters need to be tuned according to the dataset characteristics to ensure valid results. I2ACSM aims to search an attributed $pkd$-truss community that includes the query node in an Attributed and Weighted Graph (AWG) (Xie et al., 2021), ensuring that any vertex in the $k$-truss subgraph is within distance $d$ from the query node and the maximum path influence from the query node to any subgraph node exceeds $p$ (Xie et al., 2021).
TransZero-LS and TransZero-GS are proposed under TransZero, a two-phase pre-trained Graph Transformer-based community search framework that aims to find communities without ground-truth data (Wang et al., 2024b). In the offline phase, CSGphormer is pre-trained to learn node representations that encode both community information and graph topology. In the online phase, the community score is calculated by measuring the representation similarity between the query and the remaining graph nodes, and an expected score gain function is introduced to search communities with maximum expected score gain (ESG). TransZero-LS and TransZero-GS are two heuristic algorithms based on ESG for community identification.
Table 4. Representative community search algorithms.
Category Algorithm Graph Type Ground-truth
Traditional ALS (Lin et al., 2024) Undirected Temporal No
WCF-CRC (Tang et al., 2022) Dynamic No
CSD (Chen et al., 2021) Directed Multigraph  [^1] The original paper referred to the graphs as directed, but upon closer examination, we found they are more accurately characterized as directed multigraphs. No
ST-Exa (Zhang et al., 2023) Undirected No
Repeeling+ (Liao et al., 2024) Streaming Directed No
I2ACSM (Xie et al., 2021) Attributed and Weighted Yes
Learning-based TransZero-LS (Wang et al., 2024b) Undirected Yes
TransZero-GS (Wang et al., 2024b) Undirected Yes
Table 5. Parameter settings in Repeeling+ algorithm.
Dataset $|Window|$ $|Stride|/|Window|$(%) $k_{c}$ $k_{f}$
BTW17 {10k, 50k, 100k, 500k, 1M, 5M, 10M, 50M} {0.01, 0.05, 0.1, 0.5, 1, 5, 10} {0, 1} {0, 1}
Chicago_COVID {10k, 50k, 100k, 500k, 1M, 5M, 10M, 50M, 100M} {0.01, 0.05, 0.1, 0.5, 1, 5, 10} {0, 1, 2, 3} {0, 1, 2, 3}
Crawled_Dataset144 {100k, 500k, 1M, 1.5M, 2M, 2.5M} {0.001, 0.005, 0.01, 0.05, 0.1} {0, 1, 2, 3} {0, 1, 2, 3}
Crawled_Dataset26 {100k, 500k, 1M, 1.5M, 2M, 2.5M} {0.001, 0.005, 0.01, 0.05, 0.1} {0, 1, 2, 3} {0, 1, 2, 3}
Table 6. Evaluation of structural cohesiveness of CS solutions across different datasets.
Algorithm BTW17 Chicago_COVID Crawled144 Crawled26
$\mathbf{d}$ Size $Deg_{min}$ $Q_{hit}$ $\mathbf{d}$ Size $Deg_{min}$ $Q_{hit}$ $\mathbf{d}$ Size $Deg_{min}$ $Q_{hit}$ $\mathbf{d}$ Size $Deg_{min}$ $Q_{hit}$
ALS 9.24 5029.18 1.00 100 16.21 3998.33 1.18 100 15.36 9385.31 1.00 100 14.87 17043.60 1.15 100
WCF-CRC 6.88 242.80 2.19 45 10.71 280.91 8.77 52 5.64 451.59 11.12 93 9.86 287.89 3.93 39
CSD 5.67 24.00 0.67 3 4.16 19.70 59.52 10 7.70 204.22 11.14 34 6.15 125.29 8.84 20
ST-Exa 4.49 44.96 2.20 100 5.60 44.88 13.09 100 3.01 44.58 26.62 100 3.60 44.74 5.59 100
Repeeling+ 8.46 3228.53 1.49 100 6.80 998.94 18.20 100 INF INF INF INF INF INF INF INF
I2ACSM 0.55 1.61 1.14 100 1.22 3.48 101.28 100 0.97 2.41 5.95 100 0.42 1.42 4.37 100
TransZero-LS 10.00 3031.42 1.00 100 10.07 458.28 12.22 100 11.72 4536.41 1.11 100 9.10 2367.10 1.36 100
TransZero-GS - - - 0 - - - 0 - - - 0 - - - 0
6.3. The CHASE Framework
Figure 3 presents our framework coined CHASE (CoHesiveness evAluation for Social nEtworks) for evaluating CS algorithms, which includes query generation, parameter selection, graph transformation, community mapping, and cohesiveness evaluation.
We first generate $Q$ queries ($Q=100$ in our study) per dataset, each with a single query node. Queries are randomly selected from nodes in the top 50% by degree. Next, for each algorithm, datasets are transformed into algorithm-specific graph types to support its community retrieval. The varying parameter values are chosen from the original paper. For dataset-sensitive parameters, we manually test values to ensure meaningful results. For algorithms with multiple parameters, all value combinations are considered.
Using the query nodes and parameters, we obtain $Q$ communities from the transformed graph by applying the algorithm. To ensure fairness, each community is mapped back to the directed multigraphs for evaluation (i.e., cohesiveness computation). In other words, a mapped community is extracted according to nodes in the initial community and includes all edges among them. For algorithms that provide multiple solutions for a single query under various parameter combinations, the cohesiveness of these communities is averaged before comparison with other techniques.
6.4. Parameter Settings
The parameter settings for the ALS, I2ACSM, TransZero-LS, and TransZero-GS follow the original papers, with influence values and attributes for I2ACSM generated following the methods mentioned in (Xie et al., 2021). Dataset-dependent parameters for the remaining algorithms are adjusted for meaningful results. Specifically, the query time intervals for WCF-CRC and the rest of datasets are set to $t\in\{2,3\}$ and $t\in\{4,6,8,10\}$, repectively, based on the features of transformed datasets. In CSD, $k$ and $l$ are varied from 1 to 10. The size constraint $[l,h]$ in ST-Exa ranges from $[1,10]$ to $[71,80]$. The window size, stride size, $k_{c}$, and $k_{f}$ in Repeeling+ vary across datasets, thus, the parameters set for each dataset are manually tested and reported in Table 5.
We use an exponential decay function, commonly applied in fields such as biology, physics, and network analysis (Hobbie et al., 2007; Karagiannis et al., 2007; Stehlé et al., 2010), to calculate cohesiveness measures. The function is defined as $\Phi(t_{cur}-t)=e^{-\lambda(t_{cur}-t)}$, where $t_{cur}$ is the latest timestamp in each dataset. The decay rate $\lambda$ is selected from $\{0.0001,0.0005,0.001,0.005,0.01\}$, with $0.0001$ as the default. When calculating GID values, we note that the four datasets span different observation lengths and that no standard choice of $\tau_{t}$ exists. Since $W_{t_{cur}}$ uniformly scales all GID scores of communities within a dataset without affecting their relative ranking, we omit $W_{t_{cur}}$ for clarity and simplicity. As users in the four datasets lack consistent information except their ids, we used only user_id as the node feature during CSGphormer training.
Figure 3. The CHASE framework.
Figure 4. EI of communities (top row: community cohesiveness; bottom row: impact of time-decay functions).
Figure 5. SIT of communities (top row: community cohesiveness; bottom row: impact of time-decay functions).
7. Experimental Evaluation
In this section, we conduct an in-depth evaluation of the representative CS algorithms. We first assess the structural cohesiveness of their identified communities, followed by a detailed evaluation of their quality based on five psychology-informed cohesiveness measures. Learning-based algorithms were run on a Linux server with an Intel Xeon Platinum 8255C CPU (2.50GHz) and an RTX 3080 GPU (10GB memory). The remaining experiments were conducted on a Linux server with an Intel Xeon(R) Platinum 8352V CPU (2.10GHz) and 120GB memory. Recall that we do not focus on efficiency and scalability analysis here, as our goal is to understand the effectiveness of CS algorithms. The reader may refer to (Fang et al., 2020) for studies related to the former.
Figure 6. CED of communities (top row: community cohesiveness; bottom row: impact of time-decay functions).
7.1. Structural Cohesiveness
We use three commonly used evaluation metrics to assess the structural cohesiveness of CS algorithms (Fang et al., 2020; Lu et al., 2022; Xu et al., 2022): diameter ($\mathbf{d}$), size, and minimum degree ($Deg_{min}$). Here, the $Deg_{min}$ calculation includes self-loops and parallel edges. Generally, higher community quality is indicated by lower diameter and size, and higher minimum degree (Fang et al., 2020). We also report the query hit rate $Q_{hit}(\%)$, which is the ratio of queries that successfully return communities across all parameter combinations. Algorithms that fail to complete within 5 days are marked as INF and excluded from the evaluation.
Table 6 reports the results. We can make the following observations: (1) Although ALS, Repeeling+, and TransZero-LS have high $Q_{hit}$, they tend to generate larger yet sparser communities, characterized by larger $\mathbf{d}$, larger sizes, and lower $Deg_{min}$; (2) I2ACSM identifies very small communities across all datasets; (3) Compared to aforementioned algorithms, WCF-CRC, CSD, and ST-Exa are better at identifying structurally cohesive communities. However, WCF-CRC and CSD generally achieve lower hit rates, indicating difficulty in identifying target communities. Among all algorithms, ST-Exa successfully returns communities for all queries and demonstrates superior performance across all datasets, partially due to its size-constrained nature. Note that the communities (or node lists) returned by TransZero-GS are disconnected, likely due to insufficient node features during search and ignoring edge connections when selecting candidates. Hence, its results are left blank.
The poor performance of ALS may result from the smaller time scales adopted in our datasets, which hinder its ability to find nodes with high temporal proximity. However, the original paper does not discuss the time scale and its impact on CS results. Additionally, the identification of extremely large or small communities may be due to graph sparsity after multigraph transformations. The unpromising performance of the two learning-based methods may stem from the insufficient node features in our datasets, limiting CSGphormer’s ability to learn adequate features for calculating community scores. On the other hand, this highlights these methods’ stringent data requirements and their limitations when applied to datasets not specifically designed for model training.
Figure 7. GIP and GID of CS communities.
7.2. Psychology-informed Cohesiveness
Next, we report evaluation results using our psychology-informed cohesiveness measures.
Enjoyment Index (EI). The top row of Figure 4 illustrates that community cohesiveness, as indicated by EI values, differs depending on the algorithm used and is affected by the importance assigned to recent activities. For example, communities identified by Repeeling+ on the BTW show higher enjoyment when $\lambda\in\{0.0001,0.0005,0.001\}$. However, increasing the weight of recent activities reduces enjoyment due to recent negative interactions. Notably, communities identified by CSD mostly exhibit positive EI values across all datasets and are less affected by the $\lambda$.
We also examine the impact of a different time-decay function on EI, as shown in the bottom row of Figure 4. The results use a polynomial decay function, $\Phi(t-t_{cur})=(t_{cur}-t+1)^{-\mu}$ ($\mu>0$), which models a slower decline in the importance of past activities (Cormode et al., 2009). We vary $\mu$ from 0.5 to 2 in steps of 0.5. With a stronger and longer-lasting impact of recent interactions, EI values become more consistent across decay rates, and communities identified by CSD still largely maintain positive EI scores. Overall, none of the algorithms retrieved communities with consistently high EI scores.
Sentimental Interaction Tendency (SIT). SIT scores are sensitive to decay rates and the choice of decay functions. Nevertheless, the SIT results are distinguishably different from the EI results. Figure 5 shows results on all datasets. On BTW, as $\lambda$ increases, no algorithms effectively capture mutual relationships among users. Furthermore, results on C144 reveal that communities with predominantly positive EI can still include negative mutual relationships (negative SIT), and vice versa. Although CSD and Repeeling+ consider directed edges, and ALS and WCF-CRC incorporate temporal information, their communities are still inadequate to achieve higher SIT scores.
Comparative Enjoyment Degree (CED). Figure 6 plots the CED results. When $\lambda=0.0001$, some communities exhibit strong internal enjoyment, while others prefer external interactions, as indicated by the polarity of the CED scores. However, as the decay rate increases, the CED values for all communities converge towards zero, meaning the algorithms cannot find communities that increasingly prefer internal interactions over time. Similar to EI and SIT, CED is highly sensitive to decay function selection, which reflects the complexity of user interactions, as CED is calculated based on sentiments, timing, and where interactions occur. No algorithms consistently capture this complex interaction dynamics.
Group Interaction Preference (GIP). The left plot in Figure 7 reports the GIP values. ALS, WCF-CRC, and TransZero-LS tend to identify communities that are highly interactive. However, as noted in Table 6, these high GIP values may result from the large community sizes returned by these algorithms, which include most interactions. Similar patterns are observed with higher GIP values for CSD communities on the C144 and C26 datasets and lower scores for I2ACSM communities across all datasets.
Group Interaction Density (GID). The right plot in Figure 7 reports that users within communities from CSD, ST-Exa, and I2ACSM generally engage more with each other. Note that these high-GID communities also have smaller sizes and higher $Deg_{min}$.
Figure 8. Impact of sentiment analysis techniques on ATG-S measures.
7.3. Impact of Sentiment Analysis Techniques
To examine the impact of sentiment analysis techniques on evaluation and to demonstrate the broad applicability of our cohesiveness metrics, we adopt a popular lexicon-based approach, VADER (Hutto and Gilbert, 2014), to relabel the datasets. We recompute sentiment-related cohesiveness measures (i.e., EI, SIT, and CED), and compare these results with those from Llama3-labeled datasets. Figure 8 presents two results on BTW datasets. In most cases, the cohesiveness scores are largely unaffected by sentiment analysis techniques. However, when the labeling technique changes to VADER, the ST-Exa and I2ACSM communities appear more sensitive, with their values either switching polarity (e.g., ST-Exa’s EI value on BTW) or showing obvious changes (e.g., ST-Exa’s and I2ACSM’s CED values on C144).
Figure 9. Impact of parameter selection.
7.4. Impact of Parameter Selection
Next, we discuss how parameter selection affects structural and psychology-informed cohesiveness using two single-parameter algorithms, ALS and ST-Exa, on the CC dataset.
For ALS, increasing the teleportation probability $\alpha$ results in smaller communities with higher structural cohesiveness and increasing psychological cohesiveness, except GIP. Notably, GID improves significantly with larger $\alpha$, while other scores show minor increases (Figure 9, top row). For ST-Exa, members of smaller, structurally dense communities consistently achieve higher EI, CED, and GID. In contrast, members of larger communities tend to engage more, as observed in slightly higher GIP (Figure 9, bottom row).
(a) WCF-CRC
(b) ST-Exa
(c) Repeeling+
(d) I2ACSM
(e) TransZero_LS
Figure 10. Case study on the CC dataset with the query node marked in red.
Figure 11. Comparison of two types of cohesiveness.
7.5. Case Study
Finally, we present a case study using the CC dataset to compare CS results from various algorithms, evaluating cohesiveness with our proposed measures and two additional structural metrics: $k$-core and $k$-truss. The structural metrics are computed after converting the CC into an undirected graph. Here, we present the results for node “158”, as most algorithms return its communities. To simplify visualization, we only display algorithms that return communities of size 5-30. If algorithms produce multiple communities under different parameters, the largest one is selected. Besides, self-loops within the networks are not displayed, and edge widths reflect the number of interactions between users (See Figure 10).
We observe that communities returned by different algorithms vary significantly. The first three algorithms identify larger communities with more internal connections. While the communities identified by I2ACSM and TransZero_LS exhibit frequent interactions between users, their overall connectivity is sparse. For structural cohesiveness, the first three communities are 2-core and 2-truss. However, despite being truss-based, Repeeling+ and I2ACSM fail to find more prominent $k$-truss communities. I2ACSM and TransZero_LS communities only meet the 1-core requirement. In other words, existing CS techniques fail to return consistent results for the same query node.
Interestingly, our comparison of two types of cohesiveness across five communities reveals no explicit correlation between them (Figure 11). The 2-core and 2-truss communities identified by the first three algorithms lack consistent higher psychological cohesiveness. Furthermore, all methods struggle to capture psychological cohesion. That is, existing techniques return communities whose features are misaligned with the notion of cohesion in social psychology.
8. Reflections & Future Work
In this paper, we conduct the first study to extensively evaluate the cohesiveness of communities on online social networks produced by state-of-the-art CS algorithms. Our study revealed that communities identified by existing methods for the same query node lack consensus and their structural and psychological cohesiveness vary greatly, with no clear correlation between them. Despite the notion of cohesion being strongly rooted in social psychology theories, none of these generic techniques perform well on social networks w.r.t. the five psychological cohesiveness measures that are inspired by research in group cohesion. This is primarily because the existing one-size-fits-all cohesiveness measures of CS techniques do not systematically incorporate theories from social psychology in their design. This issue is further compounded by the fact that current CS techniques are either not assessed using “ground-truth” communities or rely on communities that are built without considering theories of group cohesion.
Despite these significant results, our evaluation has some limitations. Firstly, the design of our measures is constrained by sentiment analysis tools, and the simple three-class classification in our sentiment analysis may overlook nuanced sentiments inherent in specific comments. Secondly, our study is based on a single social network platform (i.e., X). However, given the fundamental limitations in the design of existing cohesiveness measures for human-interaction datasets (Section 4.2), we believe the key insights reported in this paper will also hold in other large online social networks. Thirdly, our study cannot use ground-truth communities for benchmarking because no public ground-truth datasets are created based on cohesion theories.
We believe the findings of our study can inspire a reevaluation of current community search algorithms by incorporating insights from social psychology, leading to the development of more effective solutions. Our proposed measures may be incorporated into search techniques to better identify cohesive communities in online social networks. In particular, it is paramount to create ground-truth community datasets for community search that are grounded in social psychology theories. This will not only enable the validation of our proposed psychology-informed cohesiveness measures but also assist existing CS techniques in developing more effective cohesiveness measures and algorithms for practical usage. Lastly, while this work concentrates on the community search problem, it is clear that our findings could also inspire a rethinking of various community-centric social search and analytics problems involving human interactions, such as community detection (Bedi and Sharma, 2016; Javed et al., 2018).